# latexfontFont shape

### v2

@@ Abstract
Based on the recently proposed framework of general relativistic stochastic mechanics [1, 2] and stochastic thermodynamics [3] at the ensemble level, this work focuses on general relativistic stochastic thermodynamics at the trajectory level. The first law of stochastic thermodynamics is reformulated and the fluctuation theorems are proved on this level, with emphasis on maintaining fully general covariance and on the choice of observers.
@@ 1 Introduction
One of the central problem in modern statistical physics is the origin of irreversibility in macroscopic and mesoscopic systems. This problem can be traced back to Boltzmann’s efforts in proving the second law of thermodynamics starting from deterministic mechanics, which has resulted in the well-known H-theorem [4]. However, the debates about the validity of (the assumptions of) the H-theorem have lasted for more than a century. The most acute critics to the H-theorem is reflected by Loschmidt paradox [5, 6], which roughly states that the macroscopic arrow of time cannot possibly arise from the underlying microscopic mechanics obeying time reversal symmetry (TRS). This paradox remains unresolved until the 1990s, when numerous works [7, 8, 9] emerged, revealing that the forward and reversed processes are not probabilistically equally likely, provided certain dissipative effect exists on the mechanical level. These results, known as fluctuation theorems, largely resolved the debates and paradoxes related to the H-theorem. In particular, the most questioned molecular chaos hypothesis adopted in proving the H-theorem is completely avoided in proving the fluctuation theorems.
@@
Most of the works mentioned above are carried out in the non-relativistic regime. Nowadays, it is widely acknowledged that thermodynamics and relativity are both concentrated on the universal principles which every physical system must obey. It is important to establish fluctuation theorems based on these universal principle theories. However, since the spacetime symmetry in relativity imposes stronger protection to the time reversal invariance, the extension of fluctuation theorems to the relativistic regime proves to be more difficult. The central difficulty lies in how to incorporate the breaking of time reversal invariance while still maintaining relativistic covariance. Refs. [16, 17, 18, 19, 20, 21] considered the extension of fluctuation theorems to the special relativistic regime. However, the long awaited general relativistic extension is still beyond our human’s sight.
@@
Recently, we developed a framework for dealing with stochastic mechanics on curved spacetime and investigated the general relativistic stochastic thermodynamics based on this framework [1, 2, 3]. Meanwhile, we also established a version of fluctuation theorem on an arbitrary curved Riemannian manifold [15]. The aim of the present work is to employ the framework established in [1, 2, 3] and make use of the technique introduced in [15] to construct a version of fluctuation theorem based on a fully general relativistic description of stochastic mechanics and stochastic thermodynamics.
@@
In our framework of relativistic stochastic mechanics [1, 2], it is important to liberate the observer from the coordinate system. Rather than fixing the zeroth component of the coordinate system, we utilize the observer’s proper time $t$ to to label the configuration space $\mathcal{S}_{t}$ and the space of microstates $\Sigma_{t}$, enabling our theory to possess general covariance. For this purpose, we first clarify the geometry of the space of $\mathcal{S}_{t}$ and $\Sigma_{t}$ in Sec. 2. To make the construction more self-contained, we provide a brief review of the basics of our framework of relativistic stochastic mechanics and relativistic stochastic thermodynamics in Sec. 3, and the first law of relativistic stochastic thermodynamics at trajectory level is established in this section also. The separation of the observer from the coordinate system is also important in describing the time reversal transformation (TRT) in curved spacetime. In contrast to the usual practice in special relativistic theories (including special relativistic field theories) in which the TRT is often described as a coordinate transformation $(t,x^{i})\mapsto(-t,x^{i})$, our general covariant framework calls for interpreting the TRT as a transformation from a future-directed observer to a past-directed observer. This concept is elucidated in detail in Sec. 4. Ref.[1] presents a method to establish the covariant relativistic Langevin equation from the perspective of the observer. Based on this approach, the forward and reversed processes are respectively defined as stochastic processes from the perspectives of the future-directed and past-directed observers. In Sec. 5, we demonstrate that the forward and reversed processes possess distinct probabilities, thus breaking the TRS and leading to a version of fluctuation theorem on curved spacetime manifold. Finally, in Sec. 6, we provide brief concluding remarks.
@@
We maintain the notations and conventions consistent with the previous works [1, 2]. To distinguish random variables from their realizations, the former are labeled with extra tildes. For instance, $\tilde{x}$ represents a random variable, while $x$ denotes its realization. Several manifolds of different dimensions will be relevant in our discussion. These include the $(d+1)$-dimensional spacetime manifold $\mathcal{M}$ with metric $g_{\mu\nu}(x)$ of signature (,+,,+)(-,+$\cdots$,+)( - , + , ⋯ , + ), its tangent bundle $T\mathcal{M}$ of dimension $(2d+2)$, and certain submanifolds within the tangent bundle. To distinguish tensors on these different manifolds, we introduce distinct indices. Lower-case Greek letters such as $\alpha,\beta,\mu,\nu,\rho,\ldots$, are used as concrete indices and range from 00 to $d$. Lower-case Latin letters like $i,j,k,l,m,\ldots$ are also used as concrete indices, which range from $1$ to $d$. Lastly, lower-case Latin letters $a,b,c,\ldots$ are used as abstract indices. This paper is intended to be as mathematically rigorous as possible. A more concise summary of the main results can be found in Ref.[22].
@@ 2 Geometry of the space of microstates
Statistical physics is built on top of the space of microstates. For systems consisting of classical massive particles, the space of microstates can be subdivided into the configuration space and the momentum space. In relativistic context, the configuration space is a subspace of the spacetime manifold $\mathcal{M}$ consisting of simultaneous events at a given instance of time, while the momentum space for each individual particle is a subspace of the tangent (or cotangent) space of the spacetime at a given event which obeys the mass shell condition
@@
Please be reminded that, due to the non-degeneracy of the spacetime metric $g_{\mu\nu}(x)$, the tangent and cotangent spaces are dual to each other and both are equally well for describing the momentum space. In this work, we adopt the tangent space description. When cotangent space variables appear, they are considered to be linear functions of the tangent space variables, e.g. $p_{\mu}=g_{\mu\nu}(x)p^{\nu}$.
@@
with the configuration space taken to be a subspace of the spacetime manifold $\mathcal{M}$ consisting of simultaneous events. To clarify the concept of simultaneous events, we need to introduce an arbitrary observer field which is encoded by a normalized timelike vector field $Z^{\mu}$ obeying $g_{\mu\nu}Z^{\mu}Z^{\nu}=-1$. For convenience, we shall refer to this arbitrary observer field as Alice. When considering the motion of a Brownian particle inside a heat reservoir, there is a particular observer field that comoves with the reservoir. This particular observer field will be referred to as Bob.
@@
If we consider another observer field with proper velocity $C^{\mu}=-Z^{\mu}$, referred to as Carol, whose time orientation is opposite to that of Alice, the definition of the future mass shell bundle relative to Carol will be opposite to that relative to Alice. To avoid confusion, we will designate Alice as the future-directed observer and Carol as the past-directed observer, and throughout this paper, the future and past mass shell bundles are always defined relative to Alice. This designation is arbitrary because, in the presence of TRS, the future- and the past-directed observer fields are indistinguishable on the level of deterministic mechanics.
The future mass shell bundle is also a fiber bundle based on $\mathcal{M}$. Its fiber, denoted by $(\Gamma^{+}_{m})_{x}$, is the momentum space of the relativistic particle. We use the calligraphy fonts, like $\mathcal{F}$, $\mathcal{R}$ and $\mathcal{K}$, to denote tensors on $(\Gamma_{m}^{+})_{x}$, and the cursive fonts, like $\mathscr{N}$, $\mathscr{Z}$ and $\mathscr{L}$, to denote tensors on $\Gamma_{m}^{+}$.
The future mass shell bundle still is not the space of microstates, because the base manifold $\mathcal{M}$ is not the configuration space. There are different ways to drop the temporal dimension of $\mathcal{M}$, e.g. either by fixing the zeroth component $x^{0}$ of coordinates or by fixing the proper time $t$ of Alice. The first approach lacks apparent general covariance. The core idea of our framework for relativistic stochastic mechanics is to keep the manifest general covariance and emphasize the role of observer choice. The details can be found in Ref.[1]. Here we list some of the key points:
@@
We shall also need to make use of the volume elements on $\Gamma^{+}_{m}$ and $\Sigma_{t}^{+}$. Since both of them are submanifolds of $T\mathcal{M}$, it is appropriate to begin from the geometry of $T\mathcal{M}$. The non-degenerate metric on $T\mathcal{M}$ is known as the Sasaki metric [23] $\hat{g}_{ab}$, which is determined by the metric $g_{\mu\nu}$ of the base manifold $\mathcal{M}$,
@@ 3.1 Covariant Langevin equations
Let us consider a relativistic Brownian particle carrying an electric charge $q$ and moving in a heat reservoir residing in the curved spacetime $\mathcal{M}$ and subjecting to an external electromagnetic field $F=F_{\mu\nu}{\rm d}x^{\mu}\wedge{\rm d}x^{\nu}$. A version of the corresponding general covariant Langevin equation (referred to as LEτ) that takes the proper time $\tau$ of the particle as evolution parameter reads
@@
where $\mathcal{F}^{\mu}_{\text{dp}}:=\mathcal{K}^{\mu\nu}U_{\nu}$ is the damping force with $\mathcal{K}^{\mu\nu}$ being the damping coefficient which transforms as a tensor under general coordinate transformations, $U^{\mu}$ is the proper velocity of the heat reservoir, $\mathcal{F}^{\mu}_{\text{em}}:=\frac{q}{m}F^{\mu}{}_{\nu}\tilde{p}^{\nu}_{\tau}$ is the electromagnetic force, and
@@
is the stochastic force, which is consisted of a random force term $\mathcal{R}^{\mu}{}_{\mathfrak{a}}\circ_{S}{\rm d}\tilde{w}^{\mathfrak{a}}_{\tau}/{\rm d}\tau$ encoding the Stratonovich type coupling between the stochastic amplitudes $\mathcal{R}^{\mu}{}_{\mathfrak{a}}$ with a set of independent Gaussian noises ${\rm d}\tilde{w}^{\mathfrak{a}}_{\tau}$ obeying the probability distribution
@@
in which $\nabla^{(h)}_{i}$ denotes the spatial components of the covariant derivative associated to the metric (11) on the momentum space. Both the random force and the additional stochastic force terms arise from the interaction of the Brownian particle with the heat reservoir. The additional stochastic force is required in order for the Brownian particle to be able to reach equilibrium in the long time limit [24]. Each component of $\mathcal{R}^{\mu}{}_{\mathfrak{a}}$ is assumed to be smoothly dependent on the coordinates on $\Sigma_{t}^{+}$, and for each fixed $\mathfrak{a}=1,2,\cdots d$, $\mathcal{R}^{\mu}{}_{\mathfrak{a}}$ transforms as a vector under general coordinate transformations. As did in Refs. [1, 2], we use tilded and un-tilded symbols to denote the random variables and their realizations.
Although LEτ is perfectly general covariant and encodes all necessary factors that affect the motion of the Brownian particle, there is still some drawbacks which call for an alternative version of the covariant Langevin equation. The problem is connected to the choice of evolution parameter $\tau$. Since
@@ 3.3 Thermodynamic relations
The definition of the energy of a charged relativistic particle is non-unique. For instance, both the kinematic momentum $p^{\mu}$ and canonical momentum $P^{\mu}:=p^{\mu}+qA^{\mu}$ can be used for defining the energy [25]:
@@
As an analogy of non-relativistic case, $E$ can be viewed as kinematic energy and $H$ can be viewed as the sum of kinematic energy and electromagnetic potential energy $-qZ_{\mu}A^{\mu}$. The non-uniqueness for the definition of energy also appeared in the non-relativistic stochastic thermodynamics [15]. However, such non-uniqueness does not affect the description of heat in the first law of stochastic thermodynamics. We will show that the same situation also occurs the relativistic case.
@@
The energy currents associated to the above two definitions of energy are presented as follows,
@@
where $P:=\frac{1}{d}\Delta_{\mu\nu}(U)T^{\mu\nu}[\varphi_{\text{eq}}]$ is the pressure and $n^{\mu}[\varphi]$ is particle current. Following a standard procedure as did in [29], it can be shown that $T^{\mu\nu}[\varphi_{\text{eq}}]$ and $n^{\mu}[\varphi_{\text{eq}}]$ can both be expressed in terms of the proper velocity $U^{\mu}$ of Bob: $n^{\mu}[\varphi_{\text{eq}}]$ is simply proportional to $U^{\mu}$, while $T^{\mu\nu}[\varphi_{\text{eq}}]$ takes the form of the energy-momentum tensor of a perfect fluid.
Now we would like to reinterpret eq.(49) in terms of the densities of various physical quantities as measured by Alice. This can be achieved by contracting each term in eq.(49) with $Z_{\mu}$. Recalling the fact $\beta^{\mu}=\beta U^{\mu}$, the relationship between the proper velocities of Alice and Bob, can be written as
@@
$e_{\rm A}$ is the density of the energy $E$, $h_{\rm A}$ is the density of the energy $H$, $n_{\rm A}$ is the particle number density, $s_{\rm A}:=-Z_{\mu}S^{\mu}[\varphi_{\text{eq}}]$ is the entropy density, and ${\mathcal{T}}_{\rm A}$ is the relative kinematic energy density arising from the relative motion between Alice and Bob. The subscript $\rm A$ in the notation for all these density quantities indicates that they are all defined with respect to the observer Alice. Eq. (51) can be viewed as the localized version of the Euler relation in which the coefficients of $s_{\rm A}$ and $n_{\rm A}$ should be interpreted as the temperature and chemical potential as measured by Alice. Therefore we have
@@
We can also substitute eqs. (55) and (56) into eqs. (51)-(52) to make the appearance of Euler relation simpler,
@@
It remains to introduce the trajectory entropy for the Brownian particle which plays an important role in the formulation of fluctuation theorem. Denoting the phase trajectory of the Brownian particle as $\tilde{Y}_{t}=(\tilde{y}_{t},\tilde{k}_{t})$, the trajectory entropy is defined as

### v3

@@ Abstract
Based on the recently proposed framework of general relativistic stochastic mechanics [J. Stat. Phys., 190:193, 2023; J. Stat. Phys., 190:181, 2023] and stochastic thermodynamics [SciPost Physics Core 7, 082, 2024] at the ensemble level, this work focuses on general relativistic stochastic thermodynamics at the trajectory level. The first law of stochastic thermodynamics is reformulated and the fluctuation theorems are proved on this level, with emphasis on maintaining fully general covariance and on the choice of observers.
@@ 1 Introduction
One of the central problems in modern statistical physics is the origin of irreversibility in macroscopic and mesoscopic systems. This problem can be traced back to Boltzmann’s efforts in proving the second law of thermodynamics starting from deterministic mechanics, which has resulted in the well-known H-theorem [1]. However, the debates about the validity of (the assumptions of) the H-theorem have lasted for more than a century. The most acute criticism of the H-theorem is reflected by the Loschmidt paradox [2, 3], which roughly states that the macroscopic arrow of time cannot possibly arise from the underlying microscopic mechanics obeying time reversal symmetry (TRS). This paradox remained unresolved until the 1990s, when numerous works [4, 5, 6] emerged, revealing that the forward and reversed processes are not probabilistically equally likely, provided a certain dissipative effect exists on the mechanical level. These results, known as fluctuation theorems, largely resolved the debates and paradoxes related to the H-theorem. In particular, the most questioned molecular chaos hypothesis adopted in proving the H-theorem is completely avoided in proving the fluctuation theorems.
@@
Most of the works mentioned above were carried out in the non-relativistic regime. Nowadays, it is widely acknowledged that thermodynamics and relativity are both concentrated on the universal principles that every physical system must obey. It is important to establish fluctuation theorems based on these universal principle theories. However, since the spacetime symmetry in relativity imposes stronger protection of time-reversal invariance, the extension of fluctuation theorems to the relativistic regime proves to be more difficult. The central difficulty lies in how to incorporate the breaking of time reversal invariance while still maintaining relativistic covariance. Refs. [13, 14, 15, 16, 17, 18] considered the extension of fluctuation theorems to the special relativistic regime. However, the long awaited general relativistic extension is still beyond our ability to understand.
@@
Recently, we developed a framework for dealing with stochastic mechanics on curved spacetime, and we investigated the general relativistic stochastic thermodynamics based on this framework [19, 20, 21]. Meanwhile, we also established a version of a fluctuation theorem on an arbitrary curved Riemannian manifold [12]. The aim of the present work is to employ the framework established in [19, 20, 21] and make use of the technique introduced in [12] to construct a version of a fluctuation theorem based on a fully general relativistic description of stochastic mechanics and stochastic thermodynamics.
@@
In our framework of relativistic stochastic mechanics [19, 20], it is important to liberate the observer from the coordinate system. Rather than fixing the zeroth component of the coordinate system, we utilize the observer’s proper time $t$ to label the configuration space $\mathcal{S}_{t}$ and the space of microstates $\Sigma_{t}$, enabling our theory to possess general covariance. For this purpose, we first clarify the geometry of the space of $\mathcal{S}_{t}$ and $\Sigma_{t}$ in Sec. 2. To make the construction more self-contained, we provide a brief review of the basics of our framework of relativistic stochastic mechanics and relativistic stochastic thermodynamics in Sec. 3, and the first law of relativistic stochastic thermodynamics at the trajectory level is established in this section also. The separation of the observer from the coordinate system is also important in describing the time reversal transformation (TRT) in curved spacetime. In contrast to the usual practice in special relativistic theories (including special relativistic field theories) in which the TRT is often described as a coordinate transformation $(t,x^{i})\mapsto(-t,x^{i})$, our general covariant framework calls for interpreting the TRT as a transformation from a future-directed observer to a past-directed observer. This concept is elucidated in detail in Sec. 4. Reference [19] presents a method to establish the covariant relativistic Langevin equation from the perspective of the observer. Based on this approach, the forward and reversed processes are defined, respectively, as stochastic processes from the perspectives of the future-directed and past-directed observers. In Sec. 5, we demonstrate that the forward and reversed processes possess distinct probabilities, thus breaking the TRS and leading to a version of a fluctuation theorem on a curved spacetime manifold. Finally, in Sec. 6, we provide brief concluding remarks.
@@
We maintain the notations and conventions consistent with Refs. [19, 20]. To distinguish random variables from their realizations, the former are labeled with extra tildes. For instance, $\tilde{x}$ represents a random variable, while $x$ denotes its realization. Several manifolds of different dimensions will be relevant in our discussion. These include the $(d+1)$-dimensional spacetime manifold $\mathcal{M}$ with metric $g_{\mu\nu}(x)$ of signature (,+,,+)(-,+$\cdots$,+)( - , + , ⋯ , + ), its tangent bundle $T\mathcal{M}$ of dimension $(2d+2)$, and certain submanifolds within the tangent bundle. To distinguish tensors on these different manifolds, we introduce distinct indices. Lower-case Greek letters, such as $\alpha,\beta,\mu,\nu,\rho,\ldots$, are used as concrete indices and range from 00 to $d$. Lower-case Latin letters, such as $i,j,k,l,m,\ldots$ are also used as concrete indices, which range from $1$ to $d$. Lastly, lower-case Latin letters $a,b,c,\ldots$ are used as abstract indices. This paper is intended to be as mathematically rigorous as possible. A more concise summary of the main results can be found in Ref.[22].
@@ 2 Geometry of the space of microstates
Statistical physics is built on top of the space of microstates. For systems consisting of classical massive particles, the space of microstates can be subdivided into configuration space and momentum space. In a relativistic context, configuration space is a subspace of the spacetime manifold $\mathcal{M}$ consisting of simultaneous events at a given instance of time, while the momentum space for each individual particle is a subspace of the tangent (or cotangent) space of the spacetime at a given event which obeys the mass shell condition
@@
It is important to remember that, due to the non-degeneracy of the spacetime metric $g_{\mu\nu}(x)$, the tangent and cotangent spaces are dual to each other, and both can describe the momentum space equally well. In this work, we adopt the tangent space description. When cotangent space variables appear, they are considered to be linear functions of the tangent space variables, e.g. $p_{\mu}=g_{\mu\nu}(x)p^{\nu}$.
@@
with the configuration space taken to be a subspace of the spacetime manifold $\mathcal{M}$ consisting of simultaneous events. To clarify the concept of simultaneous events, we need to introduce an arbitrary observer field that is encoded by a normalized timelike vector field $Z^{\mu}$ obeying $g_{\mu\nu}Z^{\mu}Z^{\nu}=-1$. For convenience, we shall refer to this arbitrary observer field as Alice. When considering the motion of a Brownian particle inside a heat reservoir, there is a particular observer field that comoves with the reservoir. This particular observer field will be referred to as Bob.
@@
If we consider another observer field with proper velocity $C^{\mu}=-Z^{\mu}$, referred to as Carol, whose time orientation is opposite to that of Alice, the definition of the future mass shell bundle relative to Carol will be opposite to that relative to Alice (see Fig.1). To avoid confusion, we will designate Alice as the future-directed observer and Carol as the past-directed observer, and throughout this paper, the future and past mass shell bundles are always defined relative to Alice. This designation is arbitrary because, in the presence of TRS, the future- and the past-directed observer fields are indistinguishable on the level of deterministic mechanics.
The future mass shell bundle is also a fiber bundle based on $\mathcal{M}$. Its fiber, denoted by $(\Gamma^{+}_{m})_{x}$, is the momentum space of the relativistic particle. We use calligraphy letters, such as $\mathcal{F}$, $\mathcal{R}$ and $\mathcal{K}$, to denote tensors on $(\Gamma_{m}^{+})_{x}$, and the cursive letters, such as $\mathscr{N}$, $\mathscr{Z}$ and $\mathscr{L}$, to denote tensors on $\Gamma_{m}^{+}$.
The future mass shell bundle is still not the space of microstates, because the base manifold $\mathcal{M}$ is not the configuration space. There are different ways to drop the temporal dimension of $\mathcal{M}$, e.g. either by fixing the zeroth component $x^{0}$ of coordinates or by fixing the proper time $t$ of Alice. The first approach lacks apparent general covariance. The core idea of our framework for relativistic stochastic mechanics is to keep the manifest general covariance and emphasize the role of observer choice. The details can be found in Ref.[19]. Here we list some of the key points:
@@
We shall also need to make use of the volume elements on $\Gamma^{+}_{m}$ and $\Sigma_{t}^{+}$. Since both of them are submanifolds of $T\mathcal{M}$, it is appropriate to begin from the geometry of $T\mathcal{M}$ [23]. The non-degenerate metric on $T\mathcal{M}$ is known as the Sasaki metric [24] $\hat{g}_{ab}$, which is determined by the metric $g_{\mu\nu}$ of the base manifold $\mathcal{M}$,
@@ 3.1 Covariant Langevin equations
Let us consider a relativistic Brownian particle carrying an electric charge $q$ and moving in a heat reservoir residing in the curved spacetime $\mathcal{M}$ and subjected to an external electromagnetic field $F=F_{\mu\nu}{\rm d}x^{\mu}\wedge{\rm d}x^{\nu}$. We assume that the heat reservoir has already reached thermal equilibrium, hence there is no difference between the Eckart frame and the Landau frame in defining the proper velocity $U^{\mu}$ of the reservoir. A version of the corresponding general covariant Langevin equation (referred to as LEτ) that takes the proper time $\tau$ of the particle as an evolution parameter reads
@@
where $\mathcal{F}^{\mu}_{\text{dp}}:=\mathcal{K}^{\mu\nu}U_{\nu}$ is the damping force with $\mathcal{K}^{\mu\nu}$ being the damping coefficient which transforms as a tensor under general coordinate transformations, $\mathcal{F}^{\mu}_{\text{em}}:=\frac{q}{m}F^{\mu}{}_{\nu}\tilde{p}^{\nu}_{\tau}$ is the electromagnetic force, and
@@
is the stochastic force, which consists of a random force term $\mathcal{R}^{\mu}{}_{\mathfrak{a}}\circ_{S}{\rm d}\tilde{w}^{\mathfrak{a}}_{\tau}/{\rm d}\tau$ encoding the Stratonovich type coupling between the stochastic amplitudes $\mathcal{R}^{\mu}{}_{\mathfrak{a}}$ with a set of independent Gaussian noises ${\rm d}\tilde{w}^{\mathfrak{a}}_{\tau}$ obeying the probability distribution
@@
in which $\nabla^{(h)}_{i}$ denotes the spatial components of the covariant derivative associated with the metric (11) on the momentum space. Both the random force and the additional stochastic force terms arise from the interaction of the Brownian particle with the heat reservoir. The additional stochastic force is required in order for the Brownian particle to be able to reach equilibrium in the long time limit [25]. Each component of $\mathcal{R}^{\mu}{}_{\mathfrak{a}}$ is assumed to be smoothly dependent on the coordinates on $\Sigma_{t}^{+}$, and for each fixed $\mathfrak{a}=1,2,\cdots d$, $\mathcal{R}^{\mu}{}_{\mathfrak{a}}$ transforms as a vector under general coordinate transformations. As was done in Refs. [19, 20], we use tilded and un-tilded symbols to denote the random variables and their realizations.
Although LEτ is perfectly generally covariant and encodes all necessary factors that affect the motion of the Brownian particle, there are still some drawbacks that call for an alternative version of the covariant Langevin equation. The problem is connected to the choice of evolution parameter $\tau$. Since
@@ 3.3 Thermodynamic relations
The definition of the energy of a charged relativistic particle is non-unique. For instance, both the kinematic momentum $p^{\mu}$ and the canonical momentum $P^{\mu}:=p^{\mu}+qA^{\mu}$ can be used for defining the energy [28]:
@@
As an analogy of the non-relativistic case, $E$ can be viewed as kinematic energy and $H$ can be viewed as the sum of kinematic energy and electromagnetic potential energy $-qZ_{\mu}A^{\mu}$. The non-uniqueness for the definition of energy also appeared in the non-relativistic stochastic thermodynamics [12]. However, such non-uniqueness does not affect the description of heat in the first law of stochastic thermodynamics. We will show that the same situation also occurs in the relativistic case.
@@
The energy currents associated with the above two definitions of energy are presented as follows,
@@
where $P:=\frac{1}{d}\Delta_{\mu\nu}(U)T^{\mu\nu}[\varphi_{\text{eq}}]$ is the pressure and $n^{\mu}[\varphi]$ is the particle current. Following a standard procedure as did in [27], it can be shown that $T^{\mu\nu}[\varphi_{\text{eq}}]$ and $n^{\mu}[\varphi_{\text{eq}}]$ can both be expressed in terms of the proper velocity $U^{\mu}$ of Bob: $n^{\mu}[\varphi_{\text{eq}}]$ is simply proportional to $U^{\mu}$, while $T^{\mu\nu}[\varphi_{\text{eq}}]$ takes the form of the energy-momentum tensor of a perfect fluid.
Now we would like to reinterpret eq.(49) in terms of the densities of various physical quantities as measured by Alice. This can be achieved by contracting each term in eq.(49) with $Z_{\mu}$. Recalling the fact that $\beta^{\mu}=\beta U^{\mu}$, the relationship between the proper velocities of Alice and Bob, can be written as
@@
$e_{\rm A}$ is the density of the energy $E$, $h_{\rm A}$ is the density of the energy $H$, $n_{\rm A}$ is the particle number density, $s_{\rm A}:=-Z_{\mu}S^{\mu}[\varphi_{\text{eq}}]$ is the entropy density, and ${\mathcal{T}}_{\rm A}$ is the relative kinematic energy density arising from the relative motion between Alice and Bob. The subscript $\rm A$ in the notation for all these density quantities indicates that they are all defined with respect to the observer Alice. Equation (51) can be viewed as the localized version of the Euler relation in which the coefficients of $s_{\rm A}$ and $n_{\rm A}$ should be interpreted as the temperature and chemical potential as measured by Alice. Therefore we have
@@
We can also substitute eqs. (55) and (56) into eqs. (51)-(52) to make the appearance of the Euler relation simpler,
@@
It remains to introduce the trajectory entropy for the Brownian particle, which plays an important role in the formulation of the fluctuation theorem. Denoting the phase trajectory of the Brownian particle as $\tilde{Y}_{t}=(\tilde{y}_{t},\tilde{k}_{t})$, the trajectory entropy is defined as